OpenStudy (anonymous):

Which line shows the first error in the solution? 8x-6=4x-2 (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 (3) 4x-6+6=2+6 (4) 4x=8 (5) 4x/4= 8/4 (6) x=2 (7) A. line 3 B. line 4 C. line 5 D. line 6

OpenStudy (anonymous):

give you a medal plz help me plzzzz

OpenStudy (anonymous):

i need help fast plzzz

OpenStudy (gorv):

8x-6=4x-2 subtrct 4x both side 8x-4x-6=4x-4x-2 4x-6=-2 add 6 both side 4x-6+6=-2+6 4x=4 x=1 actually it start error with line 2 there should be -2 in place of 2

OpenStudy (anonymous):

but line two is not a answer

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

okkk ...in line one -2 is there .....in line 2 how it become +2 ????? @nvisihgv

OpenStudy (anonymous):

idk @gorv

OpenStudy (anonymous):

i just know that line two is not an answer

OpenStudy (anonymous):

@Destinymasha